OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 A. y - 6 = 1/2 (x + 3) B. y = 1/2x - 6 C. y + 3 = -1/5 (x - 2) D. y - 2 = -1/5 (x + 3) (All i need is the answer)

OpenStudy (anonymous):

I'll give medal

OpenStudy (anonymous):

Ok can u share this question with people so I can get multiple answers? Thanks! ;)

OpenStudy (anonymous):

Ok thats ok, are u good at math?

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1) \] Thats the point slope formula m=slope (x1,y1)

OpenStudy (anonymous):

So is it A?

pooja195 (pooja195):

No its not

OpenStudy (anonymous):

Is it B?

pooja195 (pooja195):

Why are you guessing? Do the work.

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1) \] Thats the point slope formula m=slope (x1,y1)

OpenStudy (anonymous):

All I need is tge answer peeps

pooja195 (pooja195):

Since m=slope m=1/2 (x1,y1) x=6 y=-3

pooja195 (pooja195):

Im sure you can plug in the numbers to get your answer.

OpenStudy (anonymous):

A, B, C, or D?

pooja195 (pooja195):

Im not helping you if all you are looking for is the answer

OpenStudy (anonymous):

I am putting together a packet for a little kid, I need some help please?